f you have to pay $75.00 and have 27 bills of $1.00 and $5.00, how many bills do you need of $1.00 and $5.00 to pay.
Answer by ikleyn(52832)


Let  x  be the number of bills of  $1.
Then the number of bills of  $5  is  27-x,  in accordance with the condition.

It gives an equation
x + 5*(27-x) = 75.

To solve it,  open parentheses and reduce the like terms:

x + 5*27 - 5x = 75,
-4x = 75 - 5*27,
-4x = 75 - 135,
-4x = -60,
4x = 60,
x = 15.

Answer.  You need  15  bills of  $1  and  27-15 = 12  bills of  $5.